OpenStudy (anonymous):

can someone help, If y varies directly as x, find the constant of variation if y = 5 when x = -0.1.

OpenStudy (anonymous):

if y varies directly as x then we can write y=kx where k is constant of variation and k =y/x

OpenStudy (anonymous):

here y=5 and x= -0.1 hope u will be able to find k

OpenStudy (anonymous):

i am not sure exactly how this works in the first place

OpenStudy (anonymous):

k = -5/0.1= -50